# Multispecies inhomogeneous $t$-PushTASEP from antisymmetric fusion

### v2

@@ 1. Introduction
In this paper we focus on the version studied in [AMW24]. For a given positive integer $n$, each local state is selected from $\{0,1,\ldots,n\}$, where $1,\ldots,n$ represent the presence of one of the $n$ species of particles, and 00 corresponds to an empty site. The system evolves under a long-range stochastic dynamics on a one-dimensional periodic lattice of length $L$, with hopping rates that depend on a parameter $t$ and also on $x_{1},\ldots,x_{L}$, assigned to the lattice sites representing the inhomogeneity of the system. We refer to it as the inhomogeneous $n$-species $t$-PushTASEP, or simply PushTASEP.
Let $H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})$ denote its Markov matrix (see (10)), which appears in the continuous-time master equation. It preserves a subspace specified by the number $m_{i}$ of particles of each type $i$. Set $K_{i}=m_{0}+\cdots+m_{i-1}$. The main result of this paper, Theorem 3, is as follows:
$H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})=\frac{1}{(1-t)\prod_{i=1}^{n}(1-t^{K_{i}})}\sum_{k=0}^{n+1}(-1)^{k-1}\left.\frac{dT^{k}(z|x_{1},\ldots,x_{L})}{dz}\right|_{z=0}-\left(\sum_{j=1}^{L}\frac{1}{x_{j}}\right)\mathrm{Id},$ (1)
@@
$[T^{k}(z|x_{1},\ldots,x_{L}),T^{k^{\prime}}(z^{\prime}|x_{1},\ldots,x_{L})]=0\qquad(0\leq k,k^{\prime}\leq n+1).$ (2)
@@
A novelty here lies in the fact that $T^{k}(z|x_{1},\ldots,x_{L})$ has the auxiliary space given by the degree $k$ antisymmetric tensor representation of the quantum affine algebra $U_{t}(\widehat{sl}_{n+1})$ in a certain gauge. The corresponding quantum $R$ matrix is derived by the antisymmetric fusion, in contrast to the symmetric fusion adopted in almost all similar results obtained so far in the realm of integrable probability. [^1] It is also derived, even more simply, from the three-dimensional $L$-operator satisfying the tetrahedron equation, as reviewed in Section 3.
@@
To further expand the perspective of the result (1), let us also consider short range models, where the most extensively studied prototype is the asymmetric simple exclusion process (ASEP). Specifically, we focus on the $n$-species ASEP on the one-dimensional periodic lattice of length $L$, defined on the same state space as the aforementioned PushTASEP. The ASEP exhibits an asymmetry in the adjacent hopping rates, specified by the parameter $t$, but otherwise the system is homogeneous and possesses the $\mathbb{Z}_{L}$-translational symmetry. A variety of results have been obtained regarding the stationary states of ASEP; see, for instance, [ANP23, BW22, CDW15, CMW22, KOS24, M20] and the references therein. Let $H_{\text{ASEP}}$ denote the Markov matrix governing the continuous-time master equation (see (88a)-(88b)). It is well-known that the integrability of ASEP is attributed to the underlying commuting transfer matrices as
$H_{\text{ASEP}}=-(1-t)\frac{d}{dz}\left.\log T^{1}(z|{\bf x}={\bf 1})\right|_{z=1}-tL\,\mathrm{Id},$ (3)
@@
where $T^{1}(z|{\bf x}={\bf 1})$ is a summand corresponding to $k=1$ in (1), and ${\bf x}={\bf 1}$ indicates the specialization to the homogeneous case $x_{1}=\cdots=x_{L}=1$. This kind of origin of the “Hamiltonians” in the commuting transfer matrices is commonly referred to as Baxter’s formula (cf. [Bax82, eq. (10.14.20)]). As is customary, the evaluation is performed at the so-called “Hamiltonian point”, $z=1$ in the present setting, where $T^{1}(z|{\bf x}={\bf 1})$ reduces to a simple lattice shift operator, and the Hamiltonian becomes a sum of adjacent interaction terms under the homogeneous setting ${\bf x}={\bf 1}$.
@@
Our formula (1) is a Baxter-type formula for long-range stochastic process models, where such a Hamiltonian point does not exist due to the inherent inhomogeneity of the system. As for the second term on the right hand side, see (76) and (77) for an interpretation in terms of stationary eigenvalues. The most noteworthy feature of (1) is that it includes the superposition over all the transfer matrices corresponding to the fundamental representations of $U_{t}(\widehat{sl}_{n+1})$ for their auxiliary spaces. This is particularly intriguing because the individual transfer matrix $T^{k}(z|x_{1},\ldots,x_{L})$ is generally not stochastic; it neither satisfies non-negativity nor the so-called sum-to-unity property (cf. [KMMO16, Sec. 3.2]) in general. [^2] There are few exceptions that can be made stochastic, including the cases $k=0,1,n+1$. The alternating sum in (1) operates as an inclusion-exclusion principle, selectively extracting the allowed particle dynamics in the PushTASEP with proper transition rates, while dismissing all other unwanted channels. It would be interesting to investigate whether a similar mechanism is also effective in the generalized models where each site can accommodate more than one particle. The summation over the fundamental representations corresponds to the dimension $\binom{n+1}{0}+\binom{n+1}{1}+\cdots+\binom{n+1}{n+1}=2^{n+1}$. It indicates a further reformulation, possibly through three-dimensional integrability (cf. [K22, Chap. 18]), which is left, however, as a problem for future investigation.
@@
Let ${H_{\text{PushTASEP}}}({\bf x}={\bf 1})$ denote ${H_{\text{PushTASEP}}}(x_{1},\ldots,x_{L})$ under the homogeneous choice $x_{1}=\cdots=x_{L}=1$. This specialization presents no subtlety. The result (1) and (3) reveal that the homogeneous PushTASEP and ASEP are “sister models”, whose integrability originates from the same family of commuting transfer matrices $\{T^{k}(z|{\bf x}={\bf 1})\}$ corresponding to the fundamental representations. A direct consequence of the Yang–Baxter commutativity (2) is:
$[H_{\text{ASEP}},{H_{\text{PushTASEP}}}({\bf x}={\bf 1})]=0.$ (4)
@@
It follows that the two models share the same eigenstates. It was observed in [AMW24, Corollary 1.3] that these two models share the same stationary distribution, but this result is stronger. This property was a key motivation for the study in [AMW24], particularly in the context of stationary states. Our result provides a simple explanation for this coincidence and shows that the same stationary state is a joint eigenstate of all $T^{k}(z|x_{1},\ldots,x_{L})$. It also gives rise to an interesting question; which one among the ASEP and the homogeneous PushTASEP mixes faster, i.e. converges faster to their common stationary distribution starting from the same initial condition.
Let us comment on the inhomogeneous $n$-species $t$-PushTASEP models which are also studied from the viewpoint of vertex models in [ANP23, BW22]. Among other aspects, these models are associated with the transfer matrix whose auxiliary space corresponds to the $n+1$ dimensional vector representation of $U_{t}(\widehat{sl}_{n+1})$. In this respect, they are different from the PushTASEP considered in this paper, even though the local dynamics of pushed particles appear somewhat similar.
@@
The layout of the paper is as follows. In Section 2, we provide a precise definition of the PushTASEP following [AMW24]. In Section 3, we explain a matrix product construction of the quantum $R$ matrix $S^{k,1}(z)$ for $10\leq k\leq n+$ 1 based on the three-dimensional $L$-operator. This is a review of the results from [BS06] and [K22, Chap.11], offering a more practical approach to computing matrix elements compared to the fusion procedure detailed in Appendix A. In Section 4, we introduce the transfer matrices $T^{k}(z|x_{1},\ldots,x_{L})$ and describe their basic properties. Section 5 constitutes the core of the paper. It presents the main Theorem 3 and its proof. In Section 6, we provide further remarks on the eigenvalues of the transfer matrix and the matrix product formula for stationary states. In Section 7, we include an analogous, but much simpler known result on ASEP for reader’s convenience. Appendix A details the antisymmetric fusion. Appendix B presents another formula for $H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})$ in terms of transfer matrices associated with symmetric fusion for comparison.
@@ 2.1. Definition of $n$ species inhomogeneous $t$-PushTASEP
$\mathbb{V}({\bf m})$ $=\bigoplus_{(\sigma_{1},\ldots,\sigma_{L})\in\mathcal{S}({\bf m})}\mathbb{C}|\sigma_{1},\ldots,\sigma_{L}\rangle,$ (5)
$\mathcal{S}({\bf m})$ =$=\{(\sigma_{1},\ldots,\sigma_{L})\in\{0,\ldots,n\}^{L}\mid\delta_{i,\sigma_{1}}+\cdots+\delta_{i,\sigma_{L}}=m_{i}\,(0\leq i\leq n)\}.$ (6)
@@
$K_{i}$ $=m_{0}+\cdots+m_{i-1}\quad(0\leq i\leq n),$ (7)
$D_{\bf m}$ $=(1-t)\prod_{i=1}^{n}(1-t^{K_{i}}).$ (8)
@@
$\frac{d}{ds}|\mathbb{P}(s)\rangle={H_{\text{PushTASEP}}}(x_{1},\ldots,x_{L})|\mathbb{P}(s)\rangle$ (9)
@@
${H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle$ $=\sum_{\begin{subarray}{c}\boldsymbol{\sigma}^{\prime}\in S({\bf m})\\ \boldsymbol{\sigma}^{\prime}\neq\boldsymbol{\sigma}\end{subarray}}\sum_{j=1}^{L}\frac{1}{x_{j}}\prod_{1\leq h\leq n}w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)|\boldsymbol{\sigma}^{\prime}\rangle-\left(\sum_{j=1}^{L}\frac{[\sigma_{j}\geq 1]}{x_{j}}\right)|\boldsymbol{\sigma}\rangle,$ (10)
where we employ the Iverson bracket $[\text{true}]=1,[\text{false}]=0$ throughout. The parameter $x_{j}>0$ is associated with the lattice site $j\in\{1,\ldots,L\}$, and represents the inhomogeneity of the system at that site. The factor $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$, which constitutes the core part of ${H_{\text{PushTASEP}}}$, is a rational function of $t$ described in [AMW24, sec. 2.2]. For readers’ convenience, we recall its definition below.
Let $\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}\in S({\bf m})$. Then $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$ is defined to be zero except when the following conditions are satisfied:
@@
If case (1) holds, then $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)=1$. If case (2) holds, then let $\ell_{h}$ be the number of sites in the cyclic interval $(p(h),p^{\prime}(h))$, excluding endpoints, with value smaller than $h$ in $\boldsymbol{\sigma}$ [^3] The corresponding phrase “… smaller than $h$ in $\boldsymbol{\sigma}^{\prime}$” in [AMW24] is a misprint. Then $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$ is defined as
$w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)=\frac{(1-t)t^{\ell_{h}}}{1-t^{K_{h}}}$ (11)
using $K_{h}$ in (7). By these definitions, the first term in (10) only contains the non-diagonal terms $|\boldsymbol{\sigma}^{\prime}\rangle$ with $\boldsymbol{\sigma}^{\prime}\neq\boldsymbol{\sigma}$.
Example 1.
We consider the case $n=2$ and $L=4$. Then
$\begin{split}{H_{\text{PushTASEP}}}|0121\rangle&=\frac{|1021\rangle}{x_{2}}+\frac{|1102\rangle}{(1+t+t^{2})x_{3}}+\frac{t|2101\rangle}{(1+t+t^{2})x_{3}}+\frac{t^{2}|1201\rangle}{(1+t+t^{2})x_{3}}+\frac{|1120\rangle}{x_{4}}\\ &-\left(\frac{1}{x_{2}}+\frac{1}{x_{3}}+\frac{1}{x_{4}}\right)|0121\rangle.\end{split}$ (12)

### v3

@@ 1. Introduction
In this paper we focus on the version studied in [AMW24]. For a given positive integer $n$, each local state is selected from $\{0,1,\ldots,n\}$, where $1,\ldots,n$ represent the presence of one of the $n$ species of particles, and 0 corresponds to an empty site. The system evolves under a long-range stochastic dynamics on a one-dimensional periodic lattice of length $L$, with hopping rates that depend on a parameter $t$ and also on $x_{1},\ldots,x_{L}$, assigned to the lattice sites representing the inhomogeneity of the system. We refer to it as the inhomogeneous $n$-species $t$-PushTASEP, or simply PushTASEP.
Let $H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})$ denote its Markov matrix (see (2.6)), which appears in the continuous-time master equation. It preserves a subspace specified by the number $m_{i}$ of particles of each type $i$. Set $K_{i}=m_{0}+\cdots+m_{i-1}$. The main result of this paper, Theorem 5.1, is as follows:
$H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})=\frac{1}{(1-t)\prod_{i=1}^{n}(1-t^{K_{i}})}\sum_{k=0}^{n+1}(-1)^{k-1}\left.\frac{dT^{k}(z|x_{1},\ldots,x_{L})}{dz}\right|_{z=0}\\ -\left(\sum_{j=1}^{L}\frac{1}{x_{j}}\right)\mathrm{Id},$ (1.1)
@@
$[T^{k}(z|x_{1},\ldots,x_{L}),T^{k^{\prime}}(z^{\prime}|x_{1},\ldots,x_{L})]=0\qquad(0\leq k,k^{\prime}\leq n+1).$ (1.2)
@@
A novelty here lies in the fact that $T^{k}(z|x_{1},\ldots,x_{L})$ has the auxiliary space given by the degree $k$ antisymmetric tensor representation of the quantum affine algebra $U_{t}(\widehat{sl}_{n+1})$ in a certain gauge. The corresponding quantum $R$ matrix is derived by the antisymmetric fusion, in contrast to the symmetric fusion adopted in almost all similar results obtained so far in the realm of integrable probability. [^1] It is also derived, even more simply, from the three-dimensional $L$-operator satisfying the tetrahedron equation, as reviewed in Appendix B.
@@
To further expand the perspective of the result (1.1), let us also consider short range models, where the most extensively studied prototype is the asymmetric simple exclusion process (ASEP). Specifically, we focus on the $n$-species ASEP on the one-dimensional periodic lattice of length $L$, defined on the same state space as the aforementioned PushTASEP. The ASEP exhibits an asymmetry in the adjacent hopping rates, specified by the parameter $t$, but otherwise the system is homogeneous and possesses the $\mathbb{Z}_{L}$-translational symmetry. A variety of results have been obtained regarding the stationary states of ASEP; see, for instance, [ANP23, BW22, CDW15, CMW22, KOS24, M20, PEM09] and the references therein. Let $H_{\text{ASEP}}$ denote the Markov matrix governing the continuous-time master equation (see (7.1a)-(7.1b)). It is well-known that the integrability of ASEP is attributed to the underlying commuting transfer matrices as
$H_{\text{ASEP}}=-(1-t)\frac{d}{dz}\left.\log T^{1}(z|{\bf x}={\bf 1})\right|_{z=1}-tL\,\mathrm{Id},$ (1.3)
@@
where $T^{1}(z|{\bf x}={\bf 1})$ is a summand corresponding to $k=1$ in (1.1), and ${\bf x}={\bf 1}$ indicates the specialization to the homogeneous case $x_{1}=\cdots=x_{L}=1$. This kind of origin of the “Hamiltonians” in the commuting transfer matrices is commonly referred to as Baxter’s formula (cf. [Bax82, eq. (10.14.20)]). As is customary, the evaluation is performed at the so-called “Hamiltonian point”, $z=1$ in the present setting, where $T^{1}(z|{\bf x}={\bf 1})$ reduces to a simple lattice shift operator, and the Hamiltonian becomes a sum of adjacent interaction terms under the homogeneous setting ${\bf x}={\bf 1}$.
@@
Our formula (1.1) is a Baxter-type formula for long-range stochastic process models, where such a Hamiltonian point does not exist due to the inherent inhomogeneity of the system. The most noteworthy feature of (1.1) is that it includes the superposition over all the transfer matrices corresponding to the fundamental representations of $U_{t}(\widehat{sl}_{n+1})$ for their auxiliary spaces. This is particularly intriguing because the individual transfer matrix $T^{k}(z|x_{1},\ldots,x_{L})$ is generally not stochastic; it neither satisfies non-negativity nor the so-called sum-to-unity property (cf. [KMMO16, Sec. 3.2]) in general. [^2] There are few exceptions that can be made stochastic, including the cases $k=0,1,n+1$. The alternating sum in (1.1) operates as an inclusion-exclusion principle, selectively extracting the allowed particle dynamics in the PushTASEP with proper transition rates, while dismissing all other unwanted channels. It would be interesting to investigate whether a similar mechanism is also effective in the generalized models where each site can accommodate more than one particle. The summation over the fundamental representations corresponds to the dimension $\binom{n+1}{0}+\binom{n+1}{1}+\cdots+\binom{n+1}{n+1}=2^{n+1}$. It indicates a further reformulation, possibly through three-dimensional integrability (cf. [K22, Chap. 18]), which is left, however, as a problem for future investigation.
@@
Let ${H_{\text{PushTASEP}}}({\bf x}={\bf 1})$ denote ${H_{\text{PushTASEP}}}(x_{1},\ldots,x_{L})$ under the homogeneous choice $x_{1}=\cdots=x_{L}=1$. This specialization presents no subtlety. The result (1.1) and (1.3) reveal that the homogeneous PushTASEP and ASEP are “sister models”, whose integrability originates from the same family of commuting transfer matrices $\{T^{k}(z|{\bf x}={\bf 1})\}$ corresponding to the fundamental representations. A direct consequence of the Yang–Baxter commutativity (1.2) is:
$[H_{\text{ASEP}},{H_{\text{PushTASEP}}}({\bf x}={\bf 1})]=0.$ (1.4)
@@
It follows that the two models share the same eigenstates. It was observed in [AMW24, Corollary 1.3] that these two models share the same stationary distribution, which was a key motivation of the work. The property (1.4) provides a simple explanation for this coincidence and shows that the same stationary state is a joint eigenstate of all $T^{k}(z|x_{1},\ldots,x_{L})$. It also gives rise to an interesting question; which one among the ASEP and the homogeneous PushTASEP mixes faster, i.e. converges faster to their common stationary distribution starting from the same initial condition.
Let us comment on the inhomogeneous $n$-species $t$-PushTASEP models with arbitrary capacity, which was investigated in [BW22, ANP23]. They are formulated as integrable vertex models associated with $U_{t}(\widehat{sl}_{n+1})$. The model in [ANP23, Section 6] is a higher capacity version of our model and is a special case of a stochastic vertex model introduced in [BW22, Section 12.5]. Although the transition rates in [ANP23, Section 6] look superficially different, they turn out to be equivalent to our model at capacity $1$ because of the periodic boundary conditions. However, their results are of a very different flavor. Unlike this work, they do not give a closed form expression for the Markov matrix.
@@
The layout of the paper is as follows. In Section 2, we provide a precise definition of the PushTASEP following [AMW24]. In Section 3, we explain the antisymmetric fusion construction (cf. [KRS81]) of the quantum $R$ matrix $S^{k,1}(z)$ for $0\leq k\leq n+1$. In Section 4, we introduce the transfer matrices $T^{k}(z|x_{1},\ldots,x_{L})$ and describe their basic properties. Section 5 constitutes the core of the paper. It presents the main Theorem 5.1 and its proof. In Section 6, we provide further remarks on the eigenvalues of the transfer matrix and the matrix product formula for stationary states. In Section 7, we include an analogous, but much simpler known result on ASEP for reader’s convenience. Appendix A includes a proof of Theorem 5.1. Appendix B reviews a matrix product construction of the quantum $R$ matrix $S^{k,1}(z)$ using the three-dimensional $L$-operator, with careful adjustment of gauge and normalization. This approach, based on [BS06] and [K22, Chap. 11], provides a simpler way to compute the matrix elements than the standard fusion procedure. Appendix C presents another formula for $H_{\text{PushTASEP}}(x_{1},\ldots,x_{L})$ in terms of transfer matrices associated with the symmetric fusion for comparison.
@@ 2.1. Definition of $n$ species inhomogeneous $t$-PushTASEP
$\mathbb{V}({\bf m})$ $=\bigoplus_{(\sigma_{1},\ldots,\sigma_{L})\in\mathcal{S}({\bf m})}\mathbb{C}|\sigma_{1},\ldots,\sigma_{L}\rangle,$ (2.1)
$\mathcal{S}({\bf m})$ $=\{(\sigma_{1},\ldots,\sigma_{L})\in\{0,\ldots,n\}^{L}\mid\delta_{i,\sigma_{1}}+\cdots+\delta_{i,\sigma_{L}}=m_{i}\,(0\leq i\leq n)\}.$ (2.2)
@@
$K_{i}$ $=m_{0}+\cdots+m_{i-1}\quad(0\leq i\leq n),$ (2.3)
$D_{\bf m}$ $=(1-t)\prod_{i=1}^{n}(1-t^{K_{i}}).$ (2.4)
@@
$\frac{d}{ds}|\mathbb{P}(s)\rangle={H_{\text{PushTASEP}}}(x_{1},\ldots,x_{L})|\mathbb{P}(s)\rangle$ (2.5)
@@
${H_{\text{PushTASEP}}}|\boldsymbol{\sigma}\rangle$ $=\sum_{\begin{subarray}{c}\boldsymbol{\sigma}^{\prime}\in\mathcal{S}({\bf m})\\ \boldsymbol{\sigma}^{\prime}\neq\boldsymbol{\sigma}\end{subarray}}\sum_{j=1}^{L}\frac{1}{x_{j}}\prod_{1\leq h\leq n}w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)|\boldsymbol{\sigma}^{\prime}\rangle-\left(\sum_{j=1}^{L}\frac{[\sigma_{j}\geq 1]}{x_{j}}\right)|\boldsymbol{\sigma}\rangle,$ (2.6)
where we employ the Iverson bracket $\text{true}]=1,[\text{false}]=0$ throughout. The parameter $x_{j}>0$ is associated with the lattice site $j\in\{1,\ldots,L\}$, and represents the inhomogeneity of the system at that site. The factor $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$, which constitutes the core part of $H_{\text{PushTASEP}}$}, is a rational function of $t$ described in [AMW24, sec. 2.2]. For readers’ convenience, we recall its definition below. We first give the informal definition, followed by the formal one.
The configurations of the $t$-PushTASEP are elements in $\mathcal{S}(m)$, to be thought of as cyclic strings. The particle $b$ at site $j$ gets activated at rate $1/x_{j}$, and then starts to move clockwise, i.e. forward. When $b$ finds a particle weaker than it (i.e. with smaller value) at a site $k$ say, $b$ settles there and ‘activates’ the weaker particle with probability 1t1-t and continues to move forward with probability $t$. Since the lattice is finite with periodic boundary conditions, $b$ will find a particle to activate with probability $1$. In particular, if there are $z$ weaker particles, $b$ will displace the $i$’th one in order with probability $t^{i-1}(1-t)/(1-t^{z})$. The activated particle continues moving forward and activating particles in the same way. This ends when an activated particle settles at a vacancy. All of these moves are part of a single transition and occur instantaneously. See Figure 1 for an example.
Figure 1. All transitions going in and out of a configuration in $\mathcal{S}(1,2,1)$. The moves for the incoming transitions are shown with blue arrows.
Let $\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}\in\mathcal{S}({\bf m})$. Then $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$ is defined to be zero except when the following conditions are satisfied:
@@
If case 1 holds, then $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)=1$. If case 2 holds, then let $\ell_{h}$ be the number of sites in the cyclic interval $(p(h),p^{\prime}(h))$, excluding endpoints, with value smaller than $h$ in $\boldsymbol{\sigma}$. [^3] The corresponding phrase “… smaller than $h$ in $\boldsymbol{\sigma}^{\prime}$” in [AMW24] is a misprint. Then $w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)$ is defined as
$w^{(j)}_{\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime}}(h)=\frac{(1-t)t^{\ell_{h}}}{1-t^{K_{h}}}$ (2.7)
using $K_{h}$ in (2.3). By these definitions, the first term in (2.6) only contains the non-diagonal terms $|\boldsymbol{\sigma}^{\prime}\rangle$ with $\boldsymbol{\sigma}^{\prime}\neq\boldsymbol{\sigma}$.
Example 2.1.
We consider the example in Figure 1, which is the case of $m=(1,2,1)$, $n=2$ and $L=4$. Then
$\begin{split}{H_{\text{PushTASEP}}}|0121\rangle=&\frac{|1021\rangle}{x_{2}}+\frac{(1-t)|1102\rangle}{(1-t^{3})x_{3}}+\frac{t(1-t)|2101\rangle}{(1-t^{3})x_{3}}+\frac{t^{2}(1-t)|1201\rangle}{(1-t^{3})x_{3}}+\frac{|1120\rangle}{x_{4}}\\ &-\left(\frac{1}{x_{2}}+\frac{1}{x_{3}}+\frac{1}{x_{4}}\right)|0121\rangle.\end{split}$ (2.8)